OpenStudy (anonymous):

-6√2(-7√5-9√2) put in simplest form:

OpenStudy (solomonzelman):

\[-6\sqrt{2}(-7\sqrt{5}-9\sqrt{2})~~~⇒~~~ (-6\sqrt{2}) \times (-7\sqrt{5})~~+~~(-6\sqrt{2}) \times (-9\sqrt{2})~~~~EXPAND.\]

OpenStudy (anonymous):

the answers are A.)420+54√2 B.) 54√2+42√10 C.) 42√10+108 D.) 84

OpenStudy (solomonzelman):

\[(-6) \times (-7) \times (\sqrt{5}) \times (\sqrt{2})\]you can multiply the roots and the numbers together.\[42\sqrt{10}\]and this is plus\[(-6) \times (-9) \times (\sqrt{2}) \times (\sqrt{2})~~~~~=~~~54 \times 2~~~~~=~~~108\]\[So,~~~~~~108+42\sqrt{10}\]

OpenStudy (solomonzelman):

\[It~~~~is~~~~C\]

OpenStudy (anonymous):

okay thanks!